OpenStudy (anonymous):

Factor completely. 4x^2-2x-12

OpenStudy (anonymous):

4x^2-8x+6x-12 4x(x-2)+6(x-2) (4x+6)(x-2)

OpenStudy (anonymous):

=2(x^2-x-6) = 2(x-3)(x+2)

OpenStudy (anonymous):

4x^2-2x-12=2x^2-x-6(On dividing by 2) Or, 2x^2-x-6=0 Or, 2x^2-4x+3x-6=0 Or, 2x(x-2)+3(x-2)=0 Or, the factors are: (2x+3)(x-2)

OpenStudy (anonymous):

\[divide \ the\ eqn\ by \ 2\ \] \[2x^2 - 4x + 3x -6 =0\] \[(2x+3) (x -2) =0\] \[so, \ x= - 3/2 \ 0r\ x= 2\]

OpenStudy (anonymous):

sorry i got that wrong correct is 2(2x+3)(x-2)

OpenStudy (anonymous):

\[medal\ plz\]